OpenStudy (batman_0915):

Casey wants to wrap a gift that measures 8 inches by 6 inches by 10 inches. What is the surface area that he needs to wrap? 416 in.2 376 in.2 384 in.2 396 in.2

OpenStudy (anonymous):

Do you have an idea of the problem? If not I can work you thorugh it

OpenStudy (anonymous):

Hello? @batman_0915

OpenStudy (anonymous):

Well if you ever do come back online the problem would be worked out like this: Of the six sides: Two sides have an area 6 x 8 in^2 Two sides have an area 6 x 10 in^2 Two sides have an area 8 x 10 in^2 So the total surface area A is A = 2(6 x 8 + 6 x 10 + 8 x 10) in^2 = 2(48 + 60 + 80) in^2 = 2 * 188 in^2 = 376 in^2